OpenStudy (anonymous):

Let g(x)=3x^4-8x^3-30x^2+72x-15

OpenStudy (anonymous):

Finding g'(x) or what are we doing?

OpenStudy (anonymous):

Oh, smh sorry about that. The instructions then say: Where is g(x) increasing and where is g(x) decreasing? Find the x-values of all relative extrema.

OpenStudy (anonymous):

The first step I've done is just finding the derivative idk what's after

OpenStudy (anonymous):

Derivative tells us what in terms of this problem?

OpenStudy (anonymous):

uhhhh o.0. I don't really understande what my math teacher is doing/saying in class because he's pretty old. So you'd basically be teaching me from scratch :/

OpenStudy (anonymous):

Bummer. That's alright. The derivative gives us the slope of the original function at x. In other words, to find the slope of our original function at a point x (we'll say 3) all we have to do is plug x(3) into our derivative. What did you get for g'(x)?

OpenStudy (anonymous):

for g'(x) I found 12x^3-24x^2-60x+72

OpenStudy (anonymous):

Perfect. Did the problem give a range of values to find where g(x) is increasing or decreasing, or did it just say the whole graph?

OpenStudy (anonymous):

I can paste a link of the sample exam that I am studying from if that helps. It's nothing private

OpenStudy (anonymous):

Yeah, that's fine.:)

OpenStudy (anonymous):

Okay, so back to our problem, we need to find critical points. Critical points are found by finding the zeros of our derivative. In other words, when g'(x)=0 We have 12x^3-24x^2-60x+72=0 the next step is to factor a 12 out, giving us 12(x^3-2x^2-5x+6)=0, which simplified gives us x^3-2x^2-5x+6=0. With me so far?

OpenStudy (anonymous):

yes I found that factored "12(x^3....)" part so far

OpenStudy (anonymous):

alright, so finding the zeros of that can be a little tricky. We can use the rational root test to help with that though. Rational Root Test is given as p/q where p are your factors of your trailing constant and q is your factors of your leading coefficient. Mumbo jumbo, maybe? p are factors of 1 (coefficient of x^3) and q are factors of 6 (trailing constant) so we get \[\frac{ p }{ q }=\pm1,\pm2,\pm3,\pm6\] Now we start testing them and find that g(1)=0, which means its a root. Let's divide by (x-1) a known root now: (x^3 - 2x^2 - 5x + 6)/(x - 1) = x^2 - x - 6 Now we can factor x^2-x-6, like normal and find (x+2)(x-3) So to summarize we have (x-1)(x+2)(x-3)=0 giving us critical points at 1,-2,and 3.

OpenStudy (anonymous):

omg lol

OpenStudy (anonymous):

So now to find intervals of increasing/decreasing, we need to define those intervals based on our critical points. We end up with (-inf,-2) (-2,1) (1,3) and (3,+inf). To find if we are increasing/decreasing you pick a point in each interval and evaluate g'(x) for each point. so for our first interval (-inf,-2) let's pick 3. so g'(3)=-24. because g'(3) is negative that tells us the slope of g(x) is negative or DECREASING. Interval 2: (-2,1) x=0, g'(0)=6, so its positive and therefore g(x) is increasing from -2 to 1. Completing these steps for all 4 intervals gives us where the function is increasing and where it is decreasing...

OpenStudy (anonymous):

Ah, I'm sorry. I know it can be really overwhelming at first... let me know if i was unclear on any part.. :/ for the factoring section mathportal.org has some really helpful tools that help evaluate things and explain how to do it yourself really, really well.

OpenStudy (anonymous):

Yeah, im just a little stuck on the factoring part of the x^3...stuff

OpenStudy (anonymous):

Yeah factoring above x^2 can be really difficult, but practice really helps. Maybe try a few practice problems using this tool... http://mathportal.org/calculators/polynomials-solvers/polynomial-roots-calculator.php

OpenStudy (anonymous):

I have to run, but I'll try to get back on a little later tonight or tomorrow. Best of luck.

OpenStudy (anonymous):

ok, that is fine, thank you so much for your time however. I'll see how much more of this test I can get done

OpenStudy (anonymous):

Were you able to get through factoring and on to setting up your intervals?